# A Hybrid Systems Model of Feedback Optimization for Linear Systems: Convergence and Robustness

### v4

@@ V-C Complete Hybrid Convergence
$x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}=e^{At}x({0},{0})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau\cdot\\ B\big(u(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)\big)\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-\tilde{u}\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau B(u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-\tilde{u})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau B\tilde{u}\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau B\tilde{u}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}$. (53)
@@
$x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}=e^{At}x({0},{0})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau\cdot\\ B\big(u(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)\big)\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-\tilde{u}\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau B(u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-\tilde{u})\\ +e^{At}A^{-1}B\tilde{u}$. (56)
@@
$\left\|x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\right\|=\left\|e^{At}\right\|\left\|x({0},{0})-\tilde{x}\right\|\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}\left\|e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|\left\|u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|\left\|u(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)\right\|\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}\left\|e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|\left\|u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-\tilde{u}\right\|\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\left\|B\right\|\left\|u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-\tilde{u}\right\|$. (57)
@@
$\left\|x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\right\|=\left\|e^{At}\right\|\left\|x({0},{0})-\tilde{x}\right\|\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}\left\|e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|q^{\frac{\alpha({p-1})-1}{2}}d_{\mathcal{U}}\cdot\|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|q^{\frac{\alpha({p-1})-1}{2}}d_{\mathcal{U}}\cdot\|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}\left\|e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}\right\|d\tau\left\|B\right\|d_{\mathcal{U}}\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\left\|B\right\|d_{\mathcal{U}}.$ (58)

### v5

@@ V-C Complete Hybrid Convergence
$x(t,j)-\tilde{x}=e^{At}x({0},{0})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau\cdot\\ B\big(u(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)\big)\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-\tilde{u}\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau B(u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-\tilde{u})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau B\tilde{u}\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau B\tilde{u}-\tilde{x}$. (48)
@@
$x(t,j)-\tilde{x}=e^{At}x({0},{0})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau\cdot\\ B\big(u(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)\big)\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-\tilde{u}\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau B(u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-\tilde{u})\\ +e^{At}A^{-1}B\tilde{u}$. (51)
@@
$\left\|x(t,j)-\tilde{x}\right\|=\left\|e^{At}\right\|\left\|x({0},{0})-\tilde{x}\right\|\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}\left\|e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|\left\|u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|\left\|u(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)\right\|\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}\left\|e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|\left\|u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-\tilde{u}\right\|\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\left\|B\right\|\left\|u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-\tilde{u}\right\|$. (52)
@@
$\left\|x(t,j)-\tilde{x}\right\|=\left\|e^{At}\right\|\left\|x({0},{0})-\tilde{x}\right\|\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}\left\|e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|q^{\frac{\alpha({p-1})-1}{2}}d_{\mathcal{U}}\cdot\|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|q^{\frac{\alpha({p-1})-1}{2}}d_{\mathcal{U}}\cdot\|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}\left\|e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}\right\|d\tau\left\|B\right\|d_{\mathcal{U}}\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\left\|B\right\|d_{\mathcal{U}}.$ (53)